# Contraherent cosheaves of contramodules on Noetherian formal schemes

### 29 Mar 2026 v1

@@ 2.4. Contramodules
Part (a) is the assertion in parentheses in [31, Lemma 1.1(b)]. To prove parts (b) and (c), represent $M$ as a quotient $R$-module of a free $R$-module $F$. Then $\operatorname{Hom}_{R}(M,\mathfrak{P})$ is an $R$-submodule of the $R$-module $\operatorname{Hom}_{R}(F,\mathfrak{P})$, which an infinite product of copies of $\mathfrak{P}$. Furthermore, $\operatorname{Hom}_{R}(M,\mathfrak{P})$ is a contramodule $R$-module by Lemma 2.4.1(b). Now assertion (b) follows from the fact that the full subcategory of separated $R$-modules is closed under subobjects and infinite products in $R{\operatorname{\mathsf{--Mod}}}$. Since the full subcategory $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$ is closed under subobjects, quotient objects, and infinite products in $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}$, part (c) follows as well. For an argument based on the notion of a contramodule over a topological ring, see [16, Section B.2]. ∎
@@ 2.5. Injective torsion modules, projective and flat contramodules
The assertion follows from Lemma 2.5.1 (1) $\Leftrightarrow$ (6). For a stronger result concerning the full-and-faithfulness of the induced functor between the unbounded derived categories, see [19, Theorem 1.3]. The Noetherianity assumption can be weakened, but it cannot be simply dropped; see [24, Theorem 4.1]. ∎
For any ideal $I$ in a commutative ring $R$, the identity inclusion functor $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}\allowbreak\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ has a left adjoint functor $\Delta_{I}\colon R{\operatorname{\mathsf{--Mod}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}$[33, Examples 4.1(2–3)], [21, Example 1.3(4)]. An explicit construction of the functor $\Delta_{I}$ in the case of a finitely generated ideal $I\subset R$ can be found in [19, Proposition 2.1]; see [20, Sections 6–7] for a more detailed discussion.
@@
The 0-th left derived functor $\mathbb{L}_{0}\Lambda$ of the (neither left nor right exact) adic completion functor $\Lambda$ is the left adjoint functor $\mathbb{L}_{0}\Lambda\colon R{\operatorname{\mathsf{--Mod}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$ to the identity inclusion functor $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$[24, Proposition 1.3]. Similarly to the discussion above, for every projective $R$-module $P$, the quotseparated contramodule $R$-module $\mathbb{L}_{0}\Lambda(P)$ is projective as an object of $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$. Notice that, by the definition of the 0-th derived functor, one has $\mathbb{L}_{0}\Lambda(P)=\Lambda(P)$ for all projective $R$-modules $P$.
@@
(3) $\Longleftrightarrow$ (6) This is a particular case of a quite general lemma about an associative ring with a nilpotent ideal [16, Lemma B.10.2].
@@
(1) $\Longleftrightarrow$ (3) In the context of $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$, this also holds for any adic topological ring $R$ and does not need the Noetherianity assumption. See [17, Corollary E.1.10(a)] for an even more general result.
@@
(1), (3), or (6) $\Longrightarrow$ (7) The assumption that $R$ is Noetherian is essential here. See [16, Corollary B.8.2] or [20, Theorem 10.5] for a complete proof of (1) $\Longleftrightarrow$ (3) $\Longleftrightarrow$ (6) $\Longleftrightarrow$ (7) for Noetherian rings $R$. ∎
A discussion of the functor of contratensor product $\odot_{\mathfrak{R}}$ of discrete right $\mathfrak{R}$-modules and left $\mathfrak{R}$-contramodules, for a complete, separated topological ring $\mathfrak{R}$ where open right ideals form a base of neighborhoods of zero, can be found in [18, Section 3.3], [33, Section 5], [22, Section 2.8], [23, Sections 2.8], [35, Section 7.2], [36, Section 1]. In the context of an adic topological ring $R$ and its adic completion $\mathfrak{R}$, the natural maps
@@
are isomorphisms for any torsion $R$-module $\mathcal{M}$ and any quotseparated contramodule $R$-module $\mathfrak{P}$. This assertion, which is easy to check directly, follows also from the fact that the forgetful functor $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ is fully faithful [35, Lemma 7.11]. For this reason, we will not dwell on the contratensor product functor $\odot_{\mathfrak{R}}$, working with the tensor product $\otimes_{R}$ or $\otimes_{\mathfrak{R}}$ instead.
@@
The projective contramodule $R$-modules, i. e., the projective objects of the abelian category $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}$ are flat contramodule $R$-modules as well [31, Section 3] (see also the proof of Lemma 2.5.5 (2) $\Rightarrow$ (3) above). So flat contramodule $R$-modules need not be separated if they are not quotseparated. However, the following assertion holds.
@@
This is a special case of the much more general result of [17, Corollary E.1.7], or of the even more general [33, Corollary 6.15]. The Noetherian adic case can be found in [16, proof of Lemma B.9.2] or [20, Corollary 10.3(b)]. ∎
The following corollary is a partial generalization of [14, Lemma 3.5 and Proposition 3.6].
@@
The argument is based on the category equivalence/isomorphism $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}\simeq\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$. Part (a) is a special case of [17, Lemmas E.1.4 and E.1.5]. A further generalization can be found in [33, Corollaries 6.8 and 6.13]. Part (b) is a special case of [33, Lemma 6.10]. ∎
@@
(1) $\Longleftrightarrow$ (2) Without the Noetherianity assumption, one has to be careful. There is a natural isomorphism of the categories of quotseparated contramodules $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}\simeq\mathfrak{R}{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$, as both the categories are isomorphic to $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$. Simply put, any quotseparated contramodule $R$-module admits a unique extension of its $R$-module structure to an $\mathfrak{R}$-module structure, making it a quotseparated contramodule $\mathfrak{R}$-module; and any $R$-module morphism of quotseparated contramodule $R$-modules is an $\mathfrak{R}$-module morphism. However, a nonquotseparated contramodule $R$-module does not admit an extension of its $R$-module structure to an $\mathfrak{R}$-module structure, in general. Instead, the nonquotseparated contramodule $R$-modules have natural module structures over the ring $\Delta(R)$[21, Example 5.2(3)].
@@
(3) $\Longrightarrow$ (6) The assumption that $R$ is Noetherian is essential here. A proof can be found in [16, Lemma B.9.2] or in [20, Corollary 10.3(a)].
@@
This is [16, Theorem B.8.1]. For a stronger result concerning the full-and-faithfulness of the induced functor between the unbounded derived categories, see [19, Theorem 2.9]. The Noetherianity assumption can be weakened, but it cannot be simply dropped; see [24, Theorems 4.2 and 4.3]. ∎
@@ 2.6. Change of scalars
Let $R$ and $S$ be adic topological rings, and let $f\colon R\longrightarrow S$ be a ring homomorphism. Parts (b) and (d) of the following lemma are generalizations of [17, Lemma D.4.1(a)].
@@
Part (b): once again, for any homomorphism of commutative rings $f\colon R\longrightarrow S$ and any two ideals $I\subset R$ and $J\subset S$ such that $f(I)\subset J$, the functor $S{\operatorname{\mathsf{--Mod}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ takes $J$-contramodule $S$-modules to $I$-contramodule $R$-modules. The point is that, for any $S$-module $Q$ and any fixed element $r\in R$, there is a natural isomorphism of the Ext groups $\operatorname{Ext}^{*}_{S}(S[f(r)^{-1}],Q)\simeq\operatorname{Ext}^{*}_{R}(R[r^{-1}],Q)$, because $R[r^{-1}]$ is a flat $R$-module and $S\otimes_{R}R[r^{-1}]\simeq S[f(r)^{-1}]$. See, e. g., [34, Lemma 4.1(a)]; cf. [20, Lemma 2.1]. So the $S$-module $Q$ is an $f(r)$-contramodule if and only if the underlying $R$-module of $Q$ is an $r$-contramodule.
@@
where $I$ is an ideal of definition in $R$. As a right adjoint functor to an exact functor, the functor $f^{\diamond}$ takes injective torsion $R$-modules to injective torsion $S$-modules. The same assertions hold in the greater generality of a commutative ring homomorphism $f\colon R\longrightarrow S$ and any two ideals $I\subset R$,  $J\subset S$ such that $f(I)\subset J$[31, proof of Lemma 15.1(a)]. See [23, Section 2.9] and [32, Section 6] for a generalization to the categories of discrete modules over topological associative rings.
@@
where $I$ is an ideal of definition in $R$. As a left adjoint functor to an exact functor, the functor $f^{\#}$ takes projective contramodule $R$-modules to projective contramodule $S$-modules. Specifically, one has $f^{\#}(\Delta_{I}(R[X]))=\Delta_{J}(S[X])$ for any set $X$ (so the functor $f^{\#}$ also takes free contramodule $R$-modules to free contramodule $S$-modules). The same assertions hold in the greater generality of a commutative ring homomorphism $f\colon R\longrightarrow S$ and any two ideals $I\subset R$,  $J\subset S$ such that $f(I)\subset J$[31, proof of Lemma 15.1(b)].
@@
which can be computed by the formula $f^{\sharp}(\mathfrak{P})=\mathbb{L}_{0}\Lambda_{J}(S\otimes_{R}\mathfrak{P})$ for any $\mathfrak{P}\in R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$, where $J$ is an ideal of definition in SS[31, proof of Lemma 15.1(c)]. Moreover, for any $R$-module $P$ one has a natural isomorphism of quotseparated contramodule $S$-modules
@@
where $I$ is an ideal of definition in $R$. As a left adjoint functor to an exact functor, the functor $f^{\sharp}$ takes projective quotseparated contramodule $R$-modules to projective quotseparated contramodule $S$-modules. Specifically, one has $f^{\sharp}(\Lambda_{I}(R[X]))=\Lambda_{J}(S[X])$ for any set $X$ (so the functor $f^{\sharp}$ also takes free quotseparated contramodule $R$-modules to free quotseparated contramodule $S$-modules). See [23, Section 2.9] and [32, Section 6] for a generalization to the categories of contramodules over topological associative rings.
@@
The following corollary is a generalization of [17, Lemma D.4.3].
@@
All the assertions are based on Lemma 2.6.3(b–c). The flatness of $f^{\sharp}(\mathfrak{F})$ and $f^{\#}(\mathfrak{F})$ can be established, under the respective assumptions, using the facts that a contramodule $R$-module $\mathfrak{F}$ is flat if and only if the $R/I^{n}$-module $\mathfrak{F}/I^{n}\mathfrak{F}$ is flat for every $n\geq 1$, and similarly for contramodule $S$-modules (see Section 2.5). The formula for $f^{\sharp}(\mathfrak{F})$ follows by virtue of Lemma 2.5.6. The exactness assertion holds in view of Lemma 2.5.9(b), as the functor of projective limit of sequences of surjective maps of abelian groups is exact. Alternatively, one can also apply the argument from [17, proof of Lemma D.4.3] using Lemma 2.5.9. The result of [41, Theorem 1.2 or 2.8] or [17, Lemma E.1.3] is relevant here. ∎
@@
This is a special case of the second assertion of [32, Lemma 6.2]. The claim in question follows purely formally from the exactness assertion of Corollary 2.6.4. One needs to use the facts that the functor $f^{\sharp}$ is right exact (as a left adjoint functor), there are enough projective quotseparated contramodule $R$-modules in $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$, the projective quotseparated contramodule $R$-modules are flat (see Section 2.5), and the kernels of surjective morphisms of flat quotseparated contramodule $R$-modules are flat (see Lemma 2.5.9(a)). ∎
Now we pass from the continuous maps of adic topological rings to the tight ones. Part (b) of the following lemma is a generalization of [17, Lemma D.4.1(b)].
@@
Part (b) of the following lemma is a generalization of [17, Lemma D.4.2(a)].
@@
Similarly, for a continuous map of adic topological rings $f\colon R\longrightarrow S$ that is not tight, the functors of restriction of scalars $f_{\#}\colon S{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}$ and $f_{\sharp}\colon S{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$ do not have right adjoint functors, generally speaking. In the same example of adic topological rings $R$, $S$ and the map $f$ as above, the forgetful functor $f_{\#}=f_{\sharp}\colon S{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}$ is fully faithful, but it does not preserve infinite direct sums (because the full subcategory of $x$-contramodule modules is not closed under infinite direct sums in $k[x]{\operatorname{\mathsf{--Mod}}}$). Accordingly, the functor $f_{\#}=f_{\sharp}$ is not a left adjoint. (See the discussion in [23, Section 2.9] and the details in [32, Example 10.1].)
@@
The assertions of parts (a) and (b) are equivalent restatements of each other, because there are enough injective objecs in $S{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}$. Nevertheless, we prove them separately below. Let $I\subset R$ be an ideal of definition in $R$; then, by Lemma 2.1.6(5),  $Sf(I)\subset S$ is an ideal of definition in $S$.

### 17 Apr 2026 v2

@@ 2.4. Contramodules
Part (a) is the assertion in parentheses in [32, Lemma 1.1(b)]. To prove parts (b) and (c), represent $M$ as a quotient $R$-module of a free $R$-module $F$. Then $\operatorname{Hom}_{R}(M,\mathfrak{P})$ is an $R$-submodule of the $R$-module $\operatorname{Hom}_{R}(F,\mathfrak{P})$, which an infinite product of copies of $\mathfrak{P}$. Furthermore, $\operatorname{Hom}_{R}(M,\mathfrak{P})$ is a contramodule $R$-module by Lemma 2.4.1(b). Now assertion (b) follows from the fact that the full subcategory of separated $R$-modules is closed under subobjects and infinite products in $R{\operatorname{\mathsf{--Mod}}}$. Since the full subcategory $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$ is closed under subobjects, quotient objects, and infinite products in $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}$, part (c) follows as well. For an argument based on the notion of a contramodule over a topological ring, see [17, Section B.2]. ∎
@@ 2.5. Injective torsion modules, projective and flat contramodules
The assertion follows from Lemma 2.5.1 (1) $\Leftrightarrow$ (6). For a stronger result concerning the full-and-faithfulness of the induced functor between the unbounded derived categories, see [20, Theorem 1.3]. The Noetherianity assumption can be weakened, but it cannot be simply dropped; see [25, Theorem 4.1]. ∎
For any ideal $I$ in a commutative ring $R$, the identity inclusion functor $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}\allowbreak\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ has a left adjoint functor $\Delta_{I}\colon R{\operatorname{\mathsf{--Mod}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}$[34, Examples 4.1(2–3)], [22, Example 1.3(4)]. An explicit construction of the functor $\Delta_{I}$ in the case of a finitely generated ideal $I\subset R$ can be found in [20, Proposition 2.1]; see [21, Sections 6–7] for a more detailed discussion.
@@
The 0-th left derived functor $\mathbb{L}_{0}\Lambda$ of the (neither left nor right exact) adic completion functor $\Lambda$ is the left adjoint functor $\mathbb{L}_{0}\Lambda\colon R{\operatorname{\mathsf{--Mod}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$ to the identity inclusion functor $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$[25, Proposition 1.3]. Similarly to the discussion above, for every projective $R$-module $P$, the quotseparated contramodule $R$-module $\mathbb{L}_{0}\Lambda(P)$ is projective as an object of $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$. Notice that, by the definition of the 0-th derived functor, one has $\mathbb{L}_{0}\Lambda(P)=\Lambda(P)$ for all projective $R$-modules $P$.
@@
(3) $\Longleftrightarrow$ (6) This is a particular case of a quite general lemma about an associative ring with a nilpotent ideal [17, Lemma B.10.2].
@@
(1) $\Longleftrightarrow$ (3) In the context of $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$, this also holds for any adic topological ring $R$ and does not need the Noetherianity assumption. See [18, Corollary E.1.10(a)] for an even more general result.
@@
(1), (3), or (6) $\Longrightarrow$ (7) The assumption that $R$ is Noetherian is essential here. See [17, Corollary B.8.2] or [21, Theorem 10.5] for a complete proof of (1) $\Longleftrightarrow$ (3) $\Longleftrightarrow$ (6) $\Longleftrightarrow$ (7) for Noetherian rings $R$. ∎
A discussion of the functor of contratensor product $\odot_{\mathfrak{R}}$ of discrete right $\mathfrak{R}$-modules and left $\mathfrak{R}$-contramodules, for a complete, separated topological ring $\mathfrak{R}$ where open right ideals form a base of neighborhoods of zero, can be found in [19, Section 3.3], [34, Section 5], [23, Section 2.8], [24, Sections 2.8], [36, Section 7.2], [37, Section 1]. In the context of an adic topological ring $R$ and its adic completion $\mathfrak{R}$, the natural maps
@@
are isomorphisms for any torsion $R$-module $\mathcal{M}$ and any quotseparated contramodule $R$-module $\mathfrak{P}$. This assertion, which is easy to check directly, follows also from the fact that the forgetful functor $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ is fully faithful [36, Lemma 7.11]. For this reason, we will not dwell on the contratensor product functor $\odot_{\mathfrak{R}}$, working with the tensor product $\otimes_{R}$ or $\otimes_{\mathfrak{R}}$ instead.
@@
The projective contramodule $R$-modules, i. e., the projective objects of the abelian category $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}$ are flat contramodule $R$-modules as well [32, Section 3] (see also the proof of Lemma 2.5.5 (2) $\Rightarrow$ (3) above). So flat contramodule $R$-modules need not be separated if they are not quotseparated. However, the following assertion holds.
@@
This is a special case of the much more general result of [18, Corollary E.1.7], or of the even more general [34, Corollary 6.15]. The Noetherian adic case can be found in [17, proof of Lemma B.9.2] or [21, Corollary 10.3(b)]. ∎
The following corollary is a partial generalization of [15, Lemma 3.5 and Proposition 3.6].
@@
The argument is based on the category equivalence/isomorphism $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}\simeq\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$. Part (a) is a special case of [18, Lemmas E.1.4 and E.1.5]. A further generalization can be found in [34, Corollaries 6.8 and 6.13]. Part (b) is a special case of [34, Lemma 6.10]. ∎
@@
(1) $\Longleftrightarrow$ (2) Without the Noetherianity assumption, one has to be careful. There is a natural isomorphism of the categories of quotseparated contramodules $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}\simeq\mathfrak{R}{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$, as both the categories are isomorphic to $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$. Simply put, any quotseparated contramodule $R$-module admits a unique extension of its $R$-module structure to an $\mathfrak{R}$-module structure, making it a quotseparated contramodule $\mathfrak{R}$-module; and any $R$-module morphism of quotseparated contramodule $R$-modules is an $\mathfrak{R}$-module morphism. However, a nonquotseparated contramodule $R$-module does not admit an extension of its $R$-module structure to an $\mathfrak{R}$-module structure, in general. Instead, the nonquotseparated contramodule $R$-modules have natural module structures over the ring $\Delta(R)$[22, Example 5.2(3)].
@@
(3) $\Longrightarrow$ (6) The assumption that $R$ is Noetherian is essential here. A proof can be found in [17, Lemma B.9.2] or in [21, Corollary 10.3(a)].
@@
This is [17, Theorem B.8.1]. For a stronger result concerning the full-and-faithfulness of the induced functor between the unbounded derived categories, see [20, Theorem 2.9]. The Noetherianity assumption can be weakened, but it cannot be simply dropped; see [25, Theorems 4.2 and 4.3]. ∎
@@ 2.6. Change of scalars
Let $R$ and $S$ be adic topological rings, and let $f\colon R\longrightarrow S$ be a ring homomorphism. Parts (b) and (d) of the following lemma are generalizations of [18, Lemma D.4.1(a)].
@@
Part (b): once again, for any homomorphism of commutative rings $f\colon R\longrightarrow S$ and any two ideals $I\subset R$ and $J\subset S$ such that $f(I)\subset J$, the functor $S{\operatorname{\mathsf{--Mod}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ takes $J$-contramodule $S$-modules to $I$-contramodule $R$-modules. The point is that, for any $S$-module $Q$ and any fixed element $r\in R$, there is a natural isomorphism of the Ext groups $\operatorname{Ext}^{*}_{S}(S[f(r)^{-1}],Q)\simeq\operatorname{Ext}^{*}_{R}(R[r^{-1}],Q)$, because $R[r^{-1}]$ is a flat $R$-module and $S\otimes_{R}R[r^{-1}]\simeq S[f(r)^{-1}]$. See, e. g., [35, Lemma 4.1(a)]; cf. [21, Lemma 2.1]. So the $S$-module $Q$ is an $f(r)$-contramodule if and only if the underlying $R$-module of $Q$ is an $r$-contramodule.
@@
where $I$ is an ideal of definition in $R$. As a right adjoint functor to an exact functor, the functor $f^{\diamond}$ takes injective torsion $R$-modules to injective torsion $S$-modules. The same assertions hold in the greater generality of a commutative ring homomorphism $f\colon R\longrightarrow S$ and any two ideals $I\subset R$,  $J\subset S$ such that $f(I)\subset J$[32, proof of Lemma 15.1(a)]. See [24, Section 2.9] and [33, Section 6] for a generalization to the categories of discrete modules over topological associative rings.
@@
where $I$ is an ideal of definition in $R$. As a left adjoint functor to an exact functor, the functor $f^{\#}$ takes projective contramodule $R$-modules to projective contramodule $S$-modules. Specifically, one has $f^{\#}(\Delta_{I}(R[X]))=\Delta_{J}(S[X])$ for any set $X$ (so the functor $f^{\#}$ also takes free contramodule $R$-modules to free contramodule $S$-modules). The same assertions hold in the greater generality of a commutative ring homomorphism $f\colon R\longrightarrow S$ and any two ideals $I\subset R$,  $J\subset S$ such that $f(I)\subset J$[32, proof of Lemma 15.1(b)].
@@
which can be computed by the formula $f^{\sharp}(\mathfrak{P})=\mathbb{L}_{0}\Lambda_{J}(S\otimes_{R}\mathfrak{P})$ for any $\mathfrak{P}\in R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$, where $J$ is an ideal of definition in SS[32, proof of Lemma 15.1(c)]. Moreover, for any $R$-module $P$ one has a natural isomorphism of quotseparated contramodule $S$-modules
@@
where $I$ is an ideal of definition in $R$. As a left adjoint functor to an exact functor, the functor $f^{\sharp}$ takes projective quotseparated contramodule $R$-modules to projective quotseparated contramodule $S$-modules. Specifically, one has $f^{\sharp}(\Lambda_{I}(R[X]))=\Lambda_{J}(S[X])$ for any set $X$ (so the functor $f^{\sharp}$ also takes free quotseparated contramodule $R$-modules to free quotseparated contramodule $S$-modules). See [24, Section 2.9] and [33, Section 6] for a generalization to the categories of contramodules over topological associative rings.
@@
The following corollary is a generalization of [18, Lemma D.4.3].
@@
All the assertions are based on Lemma 2.6.3(b–c). The flatness of $f^{\sharp}(\mathfrak{F})$ and $f^{\#}(\mathfrak{F})$ can be established, under the respective assumptions, using the facts that a contramodule $R$-module $\mathfrak{F}$ is flat if and only if the $R/I^{n}$-module $\mathfrak{F}/I^{n}\mathfrak{F}$ is flat for every $n\geq 1$, and similarly for contramodule $S$-modules (see Section 2.5). The formula for $f^{\sharp}(\mathfrak{F})$ follows by virtue of Lemma 2.5.6. The exactness assertion holds in view of Lemma 2.5.9(b), as the functor of projective limit of sequences of surjective maps of abelian groups is exact. Alternatively, one can also apply the argument from [18, proof of Lemma D.4.3] using Lemma 2.5.9. The result of [42, Theorem 1.2 or 2.8] or [18, Lemma E.1.3] is relevant here. ∎
@@
This is a special case of the second assertion of [33, Lemma 6.2]. The claim in question follows purely formally from the exactness assertion of Corollary 2.6.4. One needs to use the facts that the functor $f^{\sharp}$ is right exact (as a left adjoint functor), there are enough projective quotseparated contramodule $R$-modules in $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$, the projective quotseparated contramodule $R$-modules are flat (see Section 2.5), and the kernels of surjective morphisms of flat quotseparated contramodule $R$-modules are flat (see Lemma 2.5.9(a)). ∎
Now we pass from the continuous maps of adic topological rings to the tight ones. Part (b) of the following lemma is a generalization of [18, Lemma D.4.1(b)].
@@
Part (b) of the following lemma is a generalization of [18, Lemma D.4.2(a)].
@@
Similarly, for a continuous map of adic topological rings $f\colon R\longrightarrow S$ that is not tight, the functors of restriction of scalars $f_{\#}\colon S{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}$ and $f_{\sharp}\colon S{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$ do not have right adjoint functors, generally speaking. In the same example of adic topological rings $R$, $S$ and the map $f$ as above, the forgetful functor $f_{\#}=f_{\sharp}\colon S{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}$ is fully faithful, but it does not preserve infinite direct sums (because the full subcategory of $x$-contramodule modules is not closed under infinite direct sums in $k[x]{\operatorname{\mathsf{--Mod}}}$). Accordingly, the functor $f_{\#}=f_{\sharp}$ is not a left adjoint. (See the discussion in [24, Section 2.9] and the details in [33, Example 10.1].)
@@
The assertions of parts (a) and (b) are equivalent restatements of each other, because there are enough injective objects in $S{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}$. Nevertheless, we prove them separately below. Let $I\subset R$ be an ideal of definition in $R$; then, by Lemma 2.1.6(5),  $Sf(I)\subset S$ is an ideal of definition in $S$.